OpenStudy (anonymous):

2. Choose the equation below whose axis of symmetry is x = 0. (1 point) y = x2 + 2x y = x2 − 16x + 58 y = x2 + 2 y = x2 − 4x + 2

OpenStudy (anonymous):

x=0 is the vertical line through 0 (the y-axis) So which has a vertex on the y-axis?

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

There are a few ways of knowing... one way is to know the basic graph of y = x^2 and see that y = x^2 + 2 is a vertical shift up 2 units. Or you can use the handy formula for the x-value of the vertex: x = -b/(2a) to see that y=x^2 + 2 is the only one that gives zero for this.